OpenStudy (anonymous):

Juan’s Software Service Company is a perfectly competitive firm. Juan’s total fixed cost is $25,000, its average variable cost for 1,000 service calls is $45, and its marginal revenue is $75. Juan’s makes 1,000 service calls a month. Juan’s makes an economic profit of ______ a month. The answer is $5,000, but I don't understand how.

OpenStudy (anonymous):

I understand, and I will help you :D Total fixed costs (TFC) = $25,000 Average variable costs (AVC)= $45 Marginal Revenu (MR) = $75 Total service calls = 1,000 calls (1) Find the Total costs (Total fixed costs + total variable costs) Because AVCs are just an average *$45 per each 1 call*, you need to multiply $45 by the number of total calls \[= $45 \times 1000\]Total variable costs (TVCs) = $45,000 The fixed costs are already totals, so just add them to the total variable costs and you get your total costs = $70,000 (2) Next you need to find your total revenue (TR) To do that, you need to multiply the marginal revenue (revenue per 1 call) by 1000 which is:\[ = $75 \times 1000\]Total Revenue = $75,000 (3) Total costs - total revenue = profit. \[= $75,000-70,000\]= $5,000 as profit